Work Orders in Hoover Hall

Showing all active work orders and all inactive work orders closed in the past 30 days.
WO # Requested by Description Status
1200 Benjamin Johns Hoover - Rm 2260 - Lab Chairs NO ESTIMATE REQUESTED, PROJECT WILL BE APPROVED I WOULD LIKE TO PURCHASE NEW CHAIRS FOR THIS COMPUTER LAB. Open
2506 Leigh Wiand WJ40100 - DESIGN-INT DSN-LW-2260 CHAIRS Open
97633 Eric Ohrt Order Chairs Open
97636 Eric Ohrt Install Chairs Open
99283 Michelle Grawe HOOVER-ROOM 3325 FUME HOOD COST RANGE WILL BE PROVIDED A HIGH LEVEL BALLPARK ESTIMATE IS WANTED (BASED ON INFORMATION FROM THE ABANDONED 3354 HOOVER PROJECT FROM LAST YEAR) TO ADD AN 8' COUNTER TOP HOOD TO 3342 HOOVER OR TO REPLACE THE 6' COUNTER TOP HOOD IN 3325 WITH A LONGER 15' HOOD. DESIRED COMPLETION DATE: KNOWN TIME CONSTRAINTS: WOULD LIKE AN ESTIMATE IN THE NEXT FEW WEEKS, THE WORK COULD BEGIN ANYTIME AND IDEALLY COMPLETED BY END OF MAY. DESIRED PROJECT COST: FUNDING CONTACT: CALLAHAN, MADISON Open
99284 Christine Reinders-Caron WJ25746-PR024071 - CM - NCO - CR - FUME HOODS Open
99286 Christine Reinders-Caron Order Casework Open
113877 Christine Reinders-Caron FPM FDC PM FEES Open
129475 Christine Reinders-Caron Hire DDM Electric Open
130001 Christine Reinders-Caron Hire Corn States Mechanical Open
130003 Christine Reinders-Caron Hire Bergstrom Construction Open
139125 Christine Reinders-Caron Contractor Parking Permit Open
140763 Matthew Penick From Inspection 1694 - Fume hood 081280 has a shattered sash, please replace Materials-Awaiting
142594 Christine Reinders-Caron Hire Riley Armstrong Open
158388 Christine Maduro Please route to Jeff Lamb Install additional low limit sensors for AHU3. Scott Jasper will help with the location (he has ductwork drawings) Also, install cooling coil switch at cooling coil as in other locations. On Hold
197303 Michelle Grawe We would like to investigate adding backup power to certain labs on the 3rd floor of Hoover. 3314, 3324, 3326, 3365, and possibly 3354. Open
197975 Michelle Grawe MSE would like to add the FP&M locking system to 3313, 3337, 3342, 3343, 3344, 3347, 3354 (and the door between 3344 and 3354), 3355, 3356, 3362, 3364, 3365, and possibly to 3314, 3323, and 3325, Matt Helgersen got a quote for us last fall from Baker group. We would like to move forward understanding that the costs will have to be updated for a new quote. Materials-Awaiting
198041 Michelle Grawe CoE IT and MSE would like to add FP&M locks to 2260, 2255, 2240, and 2220. CoE would pay for 2260 and 2255 and MSE would pay for 2220 and 2240, and then we would split the cost of the panel and other parts. Also, is there a way to add a noise alarm (does not need to notify anyone) to 2260 and 2255 if those doors are not latched properly? We would like to prevent the students from propping them open. Cory Johns will need to replace PG103521 with the CoE IT worktag. Open
198758 Michelle Grawe Rm 2240C - Conference Room Remodel Open
198759 Melissa Warg FDC FEES - DO NOT CHARGE TIME Open
200998 FPM Service Desk FY23 HOOVER - General Cleaning Open
202322 FPM Service Desk FY23 HOOVER - Supplies Open
206000 Ryan Braga west stairwell. HVAC not on - Not sure if this is related to the coil failures from winter, but the west stairwell HVAC system hasn't been on for some time. On Hold
224537 Abigail Zumpano Contracts Open
238305 Melissa Warg FDC MGMT FEES - DO NOT CHARGE TIME Open
240919 Jacob Wheaton The exhaust line needs to be changed from plastic piping to metal piping that goes into the fume hood in this room. Open
245662 Michael Richards Parking Permit Open
246163 Melissa Warg FDC MGMT FEES - DO NOT CHARGE TIME Open
246745 Grant Alexander Johnson I am Grant Johnson, a graduate research assistant in the MSE department. We have some equipment that requires 230V outlets. However, we donít have enough and they are in the opposite corner to where we need them. We are in 3344 Hoover Hall. We are also working on adding more equipment into the room and may require other electrical drops as well. The worktag that we would be using is: SG0500122. Let me know if there is anything else you need from me regarding this request. Thank you, Grant Johnson Assigned
251765 Eric Ohrt Mechanical Contracts Open
252657 Grant Alexander Johnson Hi, we would like to move some equipment between labs. We have a mill, lathe, load frame, laser cutter, drill press, and some other equipment that we would like to move between labs. The majority of the equipment would move from Gilman 3312 to Hoover 3344/3354. However, there would be one thing moved out of Hoover 3344 to a lab in Gilman (unknown room at this time). The worktag that we will be using is: GR-026616-00001. Let me know what questions you have and what is needed from us. Thank you, Grant Johnson Open
254289 Dustin Limoges Fan coil motor is not running and the stairwell is cold Attention Dustin Limoges Completed
258921 Nolan Christensen Problem with 2 Urinal. 1 Limed Up and the other is Leaking. Both have been valved off. Assigned
261561 Lyle Bishop Hire JCI fire to troubleshoot elevator shunt tripped instead of recalled. Open
266848 Michelle Grawe Interested in venting options for a piece of equipment in 3364 Hoover. Is there a way to add a snorkel to the south side of the room that would go through to the fume hood exhaust in 3362 Hoover? Thanks! Open
266867 Melissa Warg FDC MGMT FEES - DO NOT CHAGE TIME Open
272218 John Breon Check HVAC System, Roof Membrane is Blowing Up Assigned
272225 John Breon Estimate from Central State to replace metal cap and wall flashing on west end of top roof. Assigned
272483 Tyler Gallagher There is the ground tip from a wall plug stuck in the outlet on the upper back wall of the auditorium next to the left most double doors. Assigned
272610 Bradley Swanson Repair frozen coil on air handler 1 - AHU1 Open
272796 Michelle Grawe 3rd Floor is Too Hot Open
272940 Jordan Babler Hire Riley Armstrong Open
273184 Michelle Grawe Hi, is it possible to lower the set point for both cooling and heating for this room? The lab occupants have reported that it is warm in that lab all year round. It does feel noticeably warmer than the hallway. Thank you! Completed
273879 Kathy Anderson Broken grounding prong stuck in outlet behind Q 15 Completed
274060 Michelle Grawe Solution for high temperatures in 3365B Hoover due to equipment heat load from water chiller and XRD. Open
274062 Melissa Warg FDC MGMT FEES - DO NOT CHARGE TIME Open
279463 Makayla Lim Water is constantly leaking from above Hoover 2220BM Qi An's office space down to 22220BE Steve Martin's office. The water is leaking in front of the doors. The faculty have been made aware Assigned
279628 Michelle Grawe Entire 3rd floor is over 50% humidity. Assigned
279779 Elizabeth Salton There are 5 lights out above the projector screen. This room will be heavily utilized for orientation starting June 1st and we want all lights working. Completed
279780 Elizabeth Salton There are additional lights above the ceiling that also require replacement. This room will be heavily utilized for orientation starting June 1st and we want all lights working. Completed
280140 Ryan Braga room warm and no sound of ac system running Assigned
280176 Christine Maduro Install electric reheat coils in the lab space. Cap and remove unneeded heating water piping. * please provide a cost estimate to determine funding account Open
280192 Melissa Warg FDC MGMT FEES - DO NOT CHARGE TIME Open
280268 Eric Ohrt Install new exhaust duct per plans Attn: Grosz Assigned
281014 Abigail Zumpano Contracts Open
281192 Makayla Lim Hello, we are hoping to get a quote on a cable system similar to this picture. We have a 25 foot long wall that we would like to add this cable system to. Please let me or Andrea Mauton know what the cost would be for this. Thank you! Assigned
281283 n/a DPS Trouble On FACP Assigned
281287 Jacob Jacob Forrest We had a failure in a water line in our fume hood and have sourced a quote for a replacement part to fix it. Note from 279474 from Shane Clouser: DI faucet block is broke recommend to go through Hood manufacturer to get replacement parts DI water is off. Assigned
281501 Michael Lee Breaker tripped for the first island receptacle that sticks out of the black counter top. Tried resetting but still no power to instruments. Please send someone to resolve. Location: 3325 Hoover Hall. Phone number: 9134860069 Assigned
281516 Shane CLOUSER Call back Attn Shane Clouser Water leaking on floor Assigned
281536 Michael Lee Two new standard (16 amp) circuits need to be added in 3325 Hoover Hall. They are needed on the first island peninsula near the existing receptacle closest to the door. Phone number: 9134860069 Assigned